We have 8 groups, with 25 people in each group

Source ss df MS F
Between 8066.1 7 ? ?
Within 7565.1 192 ? ?
Total ? ? ? ?
Answer #1
Source ss df MS F
Between 8066.1 7 8066.1/7 = 1152.3 1152.3/39.4015625 = 29.2450331125828
Within 7565.1 192 7565.1/192 = 39.4015625
Total 15631.2 199